(y 2 +yz)dx+(xz+z 2 )dy+(y 2 −xy)dz=0 here P= (y^2+yz) , Q=(xz+z^2) , R=(y^2-xy)
1
Expert's answer
2020-09-03T17:32:07-0400

As per the question,

given differential equation is

("y^2+yz)dx+(xz+z^2)dy+(y^2-xy)dz=0" .....(equation 1)

Where

"y^2+yz=P" ,

"xz+z^2=Q" ,

"y^2-xy=R"

Integrate equation 1 ,

"(y^2+yz)x+c_1+(xz+z^2)y+c_2+(y^2-xy)z+c_3=0"

Where c"_1,c_2,c_3" are constant of integration

"Px+Qy+Rz+C=0"

where C="c_1+c_2+c_3"

Hence the solution of given differential is

Px+Qy+Rz+C=0